What is the concentration of a solution of sodium hydroxide, 25.0 cm3 of which requires 20.0 cm3 of hydrochloric acid of concentration 0.100mol/dm3 for neutralization?
1
Expert's answer
2011-01-28T17:00:34-0500
NaOH + HCl = NaCl + H2O
M(HCl ) = 20x10-3[dm3] * 0.1 [mol/dm3] = 20x10-4 mol
M(NaOH) = M(HCl) = 20x10-4 mol
N (NaOH) = 20x10-4 mol / 25x10-3 dm3 = 0.08 mol/dm3.